OpenStudy (anonymous):

How do you find the nth term for geometric sequences? For example, 3,6,12,24,48.

OpenStudy (cwrw238):

nth term = ar^(n-1) here a = first term = 3 and r = common ratio = 6/3 = 2 so nth term of ur sequence is 3*2^(n-1)

OpenStudy (cwrw238):

the common ratio is the number which multiplies each term to get the next eg 6 = 3*2, 12= 6*2 eyc

OpenStudy (anonymous):

Thanks so much!

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

Do you know what the formula is for arthmetic sequences?

OpenStudy (cwrw238):

nth term of an AS = a + (n-1)d a=first term, d = common difference example 2 . 4 , 6 8 .... a = 2 and d = 4-2 = 2 8th term = 2 + (8-1)*2

OpenStudy (cwrw238):

in an AS terms are of the for a, a+d, a+2d, a+3d ...